OpenStudy (bananas1234):

Simplify. (x – 3)(x^2 – 4x – 7)

OpenStudy (butterflydreamer):

To solve this question, you can re-write it as : ( x - 3 ) ( x^2 - 4x - 7) = x ( x^2 - 4x - 7 ) - 3 ( x^2 - 4x - 7) = ? Now just use the distributive property and multiply over the parentheses. Then simplify and collect like terms :)

OpenStudy (bananas1234):

ok so

OpenStudy (bananas1234):

 Like that?